A population of values has a normal distribution with μ=148.3μ=148.3 and σ=92.3σ=92.3. You intend to draw a random sample of size n=49n=49.

Find the probability that a single randomly selected value is less than 144.3.
P(X < 144.3) =

Find the probability that a sample of size n=49n=49 is randomly selected with a mean less than 144.3.
P(M < 144.3) =

Enter your answers as numbers accurate to 4 decimal places. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Let X ~ Normal ( u= 148.3, 0 = 92.3 )

n = sample size = 49

1) P( X < 144.3)

Let's find Z score as

z-- 14.3-148.3 -=-0.04334 923

So P( X < 144.3)) = P( Z < -0.04334 ) = "=NORMSDIST(-0.04334)"= 0.4827

P( M < 144.3)

Z score formula for sample mean and given value of n is as follow:

4g/m 92. 349 2 = 2-4 = 144.3 – 148.3 - = -0.30336

So P( M < 144.3)) = P( Z < -0.30336 ) = "=NORMSDIST(-0.30336)" = 0.3808